# Rewriting Pre-Training Data Boosts LLM Performance in Math and Code

### v3

@@ Abstract
The performance of large language models (LLMs) in program synthesis and mathematical reasoning is fundamentally limited by the quality of their pre-training corpora. We introduce two openly licensed datasets, released under the Llama 3.3 Community License, that significantly enhance LLM performance by systematically rewriting public data. SwallowCode ($\approx$16.1 billion tokens) refines Python snippets from The-Stack-v2 through a novel four-stage pipeline: syntax validation, pylint-based style filtering, and a two-stage LLM rewriting process that enforces style conformity and transforms snippets into self-contained, algorithmically efficient examples. Unlike prior methods that rely on exclusionary filtering or limited transformations, our transform-and-retain approach upgrades low-quality code, maximizing data utility. SwallowMath ($\approx$2.3 billion tokens) enhances Finemath-4+ by removing boilerplate, restoring context, and reformatting solutions into concise, step-by-step explanations. Within a fixed 50 billion token training budget, continual pre-training of Llama-3.1-8B with SwallowCode boosts pass@1 by +17.0 on HumanEval and +16.1 on HumanEval+ compared to Stack-Edu, surpassing the baseline model’s code generation capabilities. Similarly, substituting SwallowMath yields +12.4 accuracy on GSM8K and +7.6 on MATH. Ablation studies confirm that each pipeline stage contributes incrementally, with rewriting delivering the largest gains. All datasets, prompts, and checkpoints are publicly available, enabling reproducible research and advancing LLM pre-training for specialized domains.
@@ 1 Introduction
Figure 1: Comparison of Python-only datasets in 50 billion tokens continual pre-training setting. SwallowCode achieves the highest pass@1 on HumanEval(left) and HumanEval+(right) compared to CodeParrot-Clean, The-Stack-v1, The-Stack-v2-Smol, and Stack-Edu.
@@
Large Language Models (LLMs) have demonstrated remarkable zero-shot and few-shot capabilities across diverse tasks, yet their proficiency in mathematical reasoning and program synthesis remains constrained by the quality of pre-training corpora. Existing public datasets for specialized domains, such as The-Stack-v1 and v2 for code [1, 2] and Finemath-4+ for mathematics [3], rely primarily on rule-based extraction from web crawls (e.g., CommonCrawl) [4] or model-based scoring to filter low-quality samples. However, these approaches often retain noisy, redundant, or stylistically inconsistent data, limiting their effectiveness–particularly in the growing trend of multi-stage pretraining (or mid-training) aimed at enhancing mathematical reasoning and program synthesis (e.g., OLMo 2 [5], Nemotron-H [6], and Phi-4 [7]). For instance, as shown in Figure 1, continual pre-training of Llama-3.1-8B [8] on The-Stack-v1/v2 typically maintains baseline performance on benchmarks like HumanEval and HumanEval+ but struggles to achieve significant gains due to unaddressed data quality issues. Unlike prior methods that focus on filtering or preserving original samples, we propose rewriting pre-training corpora to eliminate noise and redundancy, delivering high-quality, self-contained data that enables efficient model learning.
This paper introduces two openly licensed datasets under the Llama 3.3 Community License, designed to advance LLM capabilities in code generation and mathematical reasoning. SwallowCode ($\approx$16.1 billion tokens) refines Python snippets from The-Stack-v2 through a novel four-stage pipeline: sequential syntax validation, pylint-based style filtering, and a two-stage LLM rewriting process that enforces style conformity and transforms snippets into algorithmically efficient, self-contained examples. By rewriting, rather than merely filtering, we eliminate noise and redundancy that persist in datasets like Stack-Edu, providing high-quality data that drives rapid accuracy improvements (Figure 1). SwallowMath ($\approx$2.3 billion tokens) transforms Finemath-4+ by removing boilerplate, restoring missing context, and reformatting solutions into concise and step-by-step explanations. Although our experiments focus on Python for ease of automated evaluation, the pipeline is language-agnostic: any programming language with a parseable syntax and a linter can benefit from the same treatment, broadening its applicability to diverse coding domains.
Continual pre-training of Llama-3.1-8B [8] for 50 billion tokens on mixed dataset with SwallowCode and multilingual corpora improves pass@1 by +17.0 on HumanEval and +16.1 on HumanEval+ compared to an equivalent budget using Stack-Edu [3]. Similarly, substituting Finemath-4+ with SwallowMath yields +12.4 accuracy on GSM8K and +7.6 on MATH. To ensure the robustness of these gains, we conducted rigorous checks for test-set leakage, streaming the entire 16.1B-token SwallowCode corpus and finding no exact matches or high-similarity documents (Jaccard similarity $\geq$ 0.8) with HumanEval or HumanEval+ prompts. All datasets, prompts, and checkpoints are publicly released, enabling reproducible research. SwallowCode and SwallowMath provide a scalable framework for enhancing LLM pre-training, driving advancements in automated reasoning and software development.
@@ 2 Related work
2.1 Open code corpora
The Stack v1 [1] aggregates approximately 3TB of permissively licensed source code from public GitHub repositories. Although deduplication and license filtering are applied, boilerplate and autogenerated files are retained without language-specific refinement. The Stack v2 [2] extends this pipeline by sourcing data from the Software Heritage archive, relaxing licensing constraints, and introducing language-specific filters to enhance data quality. With a scale of approximately 900 billion tokens, it represents a significant increase in volume over v1. However, like its predecessor, v2 adopts a filtering-only approach, preserving files without semantic enhancements, which leaves inconsistent coding styles and fragmented scripts unaddressed.
@@ 2.2 Classifier-based filtering for code corpora
Recent work has shown that classifier-based filtering strategies, such as the FineWeb-Edu approach [9], can be effective for curating high-quality datasets [10]. These methods aim to improve model performance on code-related tasks by selecting semantically rich and well-documented samples from large-scale corpora. In this context, Stack-Edu [3] represents a significant effort to create a filtered variant of StarCoder2Data [2] prioritizing high-quality code. Stack-Edu begins by selecting the 15 most prevalent programming languages from StarCoder2Data, forming a subset of approximately 450 billion tokens. To assess code quality, Stack-Edu leverages Llama-3-70B-Instruct [8] to generate synthetic annotations for 500,000 code fragments, rating each on a 0–5 scale based on educational and structural quality. These annotations train language-specific classifiers based on the StarEncoder model [11], achieving F1 scores above 0.7 for most languages when applying a threshold of 3 for binary classification. The resulting dataset comprises 125 billion tokens and demonstrates improved model convergence and higher pass@1 scores on HumanEval compared to unfiltered corpora [3]. However, Stack-Edu’s exclusionary filtering strategy discards low-scoring snippets rather than rewriting or augmenting them, leaving residual issues—such as missing context or inconsistent naming conventions in the retained data.
@@ 2.3 LLM-driven pre-training corpus rewriting
Efforts to enhance code datasets using LLMs have gained traction. Austin et al. [12] explored LLM capabilities for program synthesis, finding that human feedback significantly improves code accuracy, but their work focused on synthesis rather than dataset rewriting. Cosmopedia [13] demonstrated the potential of synthetic data generation, using Mixtral-8×7B-Instruct [14] to create high-quality text corpora, though it did not address code-specific challenges.
Jain et al. [15] applied LLM-driven code transformations to instruction-tuning exemplars, focusing on variable renaming, modularization, and comment addition. While effective for fine-tuning, their approach is limited in scope and context compared to our SwallowCode pipeline. Their transformations address only a subset of stylistic improvements, whereas our Style-Guided Code Rewriting (SGCR) pipeline comprehensively enforces ten criteria from the Google Python Style Guide, including descriptive variable naming, effective type annotations, modular function design, error handling, and readability-focused formatting (see Section 3.3.1). Furthermore, our Self-Contained Optimization Rewriting (SCOR) pipeline introduces semantic enhancements—ensuring self-containment by resolving dependencies, optimizing algorithms, and transforming trivial snippets into instructive examples—which are absent in Jain et al.’s work (see Section 3.3.2). Additionally, Jain et al. target instruction-tuning data, a smaller and more curated dataset, whereas SwallowCode systematically rewrites large-scale pre-training corpora, a more challenging task due to the diversity and volume of code samples. Our pipeline integrates rigorous preprocessing with syntax error and linter-based filtering to ensure high-quality inputs for rewriting.
SwallowCode’s transform-and-retain paradigm upgrades low-quality code rather than discarding it, addressing limitations of existing datasets like The Stack v1 [1] and v2 [2], and Stack-Edu [3], which rely on filtering. By combining filtering, SGCR, and SCOR, SwallowCode produces a high-quality corpus that significantly improves performance on HumanEval and HumanEval+ benchmarks, advancing the state-of-the-art in code dataset curation.
@@ 2.4 Synthetic Data Generation for Code
Recent approaches, such as Magpie [16], leverage high-performing LLMs to generate synthetic instruction-tuning datasets from prompts and desired characteristics without relying on existing data. However, we opted against generating synthetic code data from scratch for SwallowCode due to two key limitations. First, previous work [17] demonstrates that low-diversity synthetic data restricts LLM performance. Second, achieving high diversity in synthetic code datasets requires diverse topics and keywords as seeds, as seen in Nemotron-4 340B’s synthetic instruction data [18]. For code, defining such seeds (e.g., varied algorithmic paradigms or problem domains) remains an open challenge, as no established methodology ensures sufficient diversity across programming tasks. Instead, our approach leverages high-quality code from The-Stack-v2, filtered for syntactic and stylistic rigor (Section 3.2), and applies LLM-driven rewriting to enhance quality while preserving the inherent diversity of real-world code. This transform-and-retain strategy maximizes data utility and avoids the risks of synthetic data homogeneity.
@@ 3 Construction of the code corpus
The development of SwallowCode is driven by an empirical and exploratory approach, informed by data ablation experiments evaluating each stage of the data processing pipeline, as illustrated in Figure 2. Specifically, we continually pre-trained the baseline model with the dataset only differing code text subset: before and after applying specific stage in the pipeline, and decide if we should adopt or discard the stage based on the evaluation results. In this section, we present the experimental results and describe the design choices shaping the pipeline. Please refer to Appendix H.1 for detailed results. To ensure full reproducibility, all code, configurations, and supporting materials are publicly available at https://github.com/rioyokotalab/swallow-code-math.
@@ 3.1 Experimental setup
To evaluate the impact of each design choice in our data processing pipeline, we conduct systematic data ablation studies. Each ablation trains a model that differs only in the target pre-training dataset, holding all other factors constant, including model architecture, parameter count, non-target data, total token budget, and hyperparameters. Specifically, we perform continual pre-training starting from Llama-3.1-8B, using a total of approximately 50 billion tokens, a maximum sequence length of 8,192, a global batch size of approximately 4 million tokens, and the Llama-3 tokenizer. The target dataset is processed with less than one epoch within each ablation. We evaluate model checkpoints approximately every 10 billion tokens using ten downstream benchmarks.
Continual pre-training is conducted using Megatron-LM (version core_r0.9.0) [19]. For evaluation, we utilize the BigCodeBench [20] and lm-evaluation-harness [21] on the following benchmarks: OpenBookQA [22], TriviaQA [23], HellaSwag [24], SQuAD 2.0 [25], XWINO [26], MMLU [27], GSM8K [28], BBH [29], HumanEval [30], and HumanEval+ [31]. The effectiveness of the code corpus is specifically assessed using HumanEval and HumanEval+. The pre-training data mixture consists of 84% multilingual text and 16% code. Detailed proportions and data source are provided in Appendix A.4.1, with comprehensive training hyperparameters reported in Appendix A.1. All ablation models, associated checkpoints, and supporting materials are publicly available in our Hugging Face repository to ensure full reproducibility.
@@ 3.2 Filtering
To construct the SwallowCode corpus, a high-quality Python code dataset, we implement a rigorous filtering pipeline starting with the-stack-v2-train-smol-ids [2] as the base dataset. The filtering process is critical to ensure that only syntactically correct and well-structured code is retained, enhancing downstream performance in code generation tasks. We focus exclusively on Python code to maintain consistency across ablation studies and enable fair comparisons with existing public corpora. Our pipeline employs two key filtering techniques, syntax error filtering and linter-based filtering, that significantly improve code quality. We also evaluated LLM-based scoring in ablation experiments but did not adopt it in the final pipeline due to its limited performance gains relative to computational cost. Figure 3 summarizes the performance of these methods on the HumanEval and HumanEval+ benchmarks, demonstrating the impact of our filtering strategy.
@@ 3.2.1 Syntax error filtering
Despite the heuristic filtering in the BigCode project, the-stack-v2-train-smol-ids [2] includes Python code samples with invalid syntax according to Python 3.10 specifications. To address this, we apply syntax error filtering by compiling each code sample using Python’s built-in compile() function, discarding any samples that fail to compile. This process reduces the dataset from approximately 41 million to 37 million samples, a 9.7% reduction. As shown in Figure 3, removing syntactically invalid samples improves performance on HumanEval and HumanEval+  [^1] The performance trajectory of The Stack v2 exhibits an initial decline followed by a recovery, consistent with forgetting and subsequent adaptation observed in prior continual pretraining studies [32]. This behavior is not particularly noteworthy.. Consequently, we adopt syntax-error filtering as a standard step in all subsequent experiments.
@@ 3.2.2 Linter-based filtering
Beyond syntactic correctness, code quality depends on adherence to coding standards. Many samples in the initial dataset exhibit poor structure, generating numerous warnings when analyzed by static analysis tools. We employ Pylint, a widely-used Python linter, to enforce a quality threshold, excluding samples with scores below 7.0 on a 0–10 scale based on rule violations. Additionally, we penalize overly verbose comments using a custom heuristic scoring algorithm (detailed in Appendix B). This step reduces the dataset from 36.7 million to 24.1 million samples, a 34.3% reduction. Figure 3 illustrates the performance gains of more than 1 point in HumanEval and HumanEval + achieved by linter-based filtering. Consequently, we adopt linter-based filtering, with a threshold of 7.0, as a standard step in all subsequent experiments.
3.2.3 LLM-based scoring filtering
Recent approaches leverage large language models (LLMs) to generate synthetic annotations for training quality classifiers, which are then used to filter web-scale corpora by retaining high-quality samples (e.g., FineWeb [9], Stack-Edu [3]). Instead of training a separate classifier, we directly prompt Llama-3.3-70B-Instruct [8] to evaluate each Python code snippet on a scale of 0–10, based on ten criteria, including code readability, modularity, and adherence to naming conventions, derived from the Google Python Style Guide. The detailed scoring prompt and the distribution of the quality scores are provided in Appendix C.

### v4

# arXiv:2505.02881v4[cs.LG] 01 Mar 2026

@@ Abstract
The performance of large language models (LLMs) in program synthesis and mathematical reasoning is fundamentally limited by the quality of their pre-training corpora. We introduce two openly licensed pre-training datasets, released under the Llama 3.3 Community License, that significantly enhance LLM performance by systematically rewriting public data. SwallowCode ($\approx$16.1 billion tokens) refines Python snippets from The-Stack-v2 through a novel four-stage pipeline: syntax validation, pylint-based style filtering, and a two-stage LLM rewriting process that enforces style conformity and transforms snippets into self-contained, algorithmically efficient examples. Unlike prior methods that rely on exclusionary filtering or limited transformations, our transform-and-retain approach refines low-quality code, maximizing data utility. SwallowMath ($\approx$2.3 billion tokens) enhances Finemath-4+ by removing boilerplate, restoring context, and reformatting solutions into concise, step-by-step explanations. Within a fixed 50 billion token training budget, continual pre-training of Llama-3.1-8B with SwallowCode boosts pass@1 by +17.0 on HumanEval and +16.1 on HumanEval+ compared to Stack-Edu, surpassing the baseline model’s code generation capabilities. Similarly, substituting SwallowMath yields +12.4 accuracy on GSM8K and +7.6 on MATH. Ablation studies confirm that each pipeline stage contributes incrementally, with rewriting yielding the largest gains. By releasing datasets, prompts, checkpoints, and pipeline code, we ensure reproducibility and provide a transferable transform-and-retain methodology that can be adapted to other base models and LLM rewriting setups.
@@ 1 Introduction
Figure 1: Comparison of Python-only datasets in a 50 billion tokens continual pre-training setting. SwallowCode achieves the highest pass@1 on HumanEval(left) and HumanEval+(right) compared to CodeParrot-Clean, The-Stack-v1, The-Stack-v2-Smol, and Stack-Edu.
@@
Large Language Models (LLMs) have demonstrated remarkable zero-shot and few-shot capabilities across diverse tasks, yet their proficiency in mathematical reasoning and program synthesis remains constrained by the quality of pre-training corpora. Existing public datasets for specialized domains, such as The-Stack-v1 and v2 for code (Kocetkov et al., 2022; Lozhkov et al., 2024) and Finemath-4+ for mathematics (Allal et al., 2025), rely primarily on rule-based extraction from web crawls (e.g., CommonCrawl) (Paster et al., 2023) or model-based scoring to filter low-quality samples. However, these approaches often retain noisy, redundant, or stylistically inconsistent data, limiting their effectiveness–particularly in the growing trend of multi-stage pre-training (or mid-training) aimed at enhancing mathematical reasoning and program synthesis (e.g., OLMo 2 (OLMo et al., 2025), Nemotron-H (NVIDIA et al., 2025), and Phi-4 (Abdin et al., 2024)).
[…]
We deliberately avoided continual pre-training from near-frontier models (e.g., Qwen-2.5/3, gpt-oss). These models already achieve very high performance in mathematics and code, so additional data yields only marginal gains. Such settings obscure whether improvements stem from the dataset itself or from incidental model factors, confounding the attribution of results. In contrast, Llama-3.1-8B (Grattafiori et al., 2024) provides a more informative starting point: it is strong enough to serve as a realistic baseline for open-community efforts, yet remains sufficiently below the frontier to avoid saturation, thereby allowing improvements to be clearly attributed to dataset quality.
Under this protocol, continual pre-training for 50B tokens on a mixed dataset containing SwallowCode and multilingual corpora improves pass@1 by +17.0 on HumanEval and +16.1 on HumanEval+ compared to an equivalent budget using Stack-Edu. Substituting Finemath-4+ with SwallowMath yields +12.4 accuracy on GSM8K and +7.6 on MATH. To verify generality beyond the Llama-3 family, we also conduct 20B-token continual pre-training from Qwen2-7B (Yang et al., 2024), where the model trained with SwallowCode surpasses Stack-Edu, improving HumanEval by +10.3 and HumanEval+ by +10.3 (see Appendix B). We conduct rigorous decontamination checks for SwallowCode against HumanEval and HumanEval+ prompts (no exact matches or high-similarity documents; see Appendix H) and apply analogous procedures for SwallowMath against GSM8K and MATH (no contamination). We further examine potential self-contamination from the rewriting LLM (Llama-3.3-70B-Instruct) using post-cutoff benchmarks (Appendix H). All datasets, prompts, and checkpoints are publicly released to ensure reproducibility and to enable adaptation to other base models and rewriting using LLMs.
@@ 2.1 Classifier-based filtering for code corpora
[…]
@@ 2.2 LLM-driven pre-training corpus rewriting
Efforts to improve datasets using LLMs have gained traction. Cosmopedia (Ben Allal et al., 2024) demonstrated the potential of synthetic data generation, using Mixtral-8×7B-Instruct (Jiang et al., 2024) to create high-quality text corpora, although it did not address code-specific challenges. Related techniques have also emerged in other domains, particularly for general web data refinement. For example, Rephrasing the Web (Maini et al., 2024) and Nemotron-CC (Su et al., 2025) transform noisy Common Crawl text into improved corpora through large-scale rephrasing or QA-style conversions. These approaches, however, target non-code/math domains and often perform format changes that overlap with instruction-tuning objectives. In contrast, our method applies domain-specific rewriting while ensuring that the outputs remain pure code snippets, thereby preserving the original data’s semantics. Unlike web data rephrasing approaches that introduce style conversions and may implicitly boost instruction-following abilities during pre-training, our approach is not designed to gain such capabilities but rather to improve the quality of pre-training data itself. This distinction makes our work fundamentally different from prior web data rephrasing efforts and highlights its unique contribution. MegaMath (Zhou et al., 2025) represents another recent attempt to refine math data, but adopts a QA-style interleaved text-code format. This design aligns more closely with general web data refinement methods that restructure content into instruction-response format, in contrast to our approach of retaining pure, self-contained snippets.
Jain et al. (2023) is closest to our work, applying LLM-driven code transformations to instruction tuning exemplars, e.g., variable renaming, modularization, and comment addition. While effective for fine-tuning, their approach is limited in scope and context compared to our SwallowCode pipeline. Their transformations address only a subset of stylistic improvements, whereas our Style-Guided Code Rewriting (SGCR) pipeline comprehensively enforces ten criteria from the Google Python Style Guide, including descriptive variable naming, effective type annotations, modular function design, error handling, and readability-focused formatting (see Section 3.3.1). Furthermore, our Self-Contained Optimization Rewriting (SCOR) pipeline introduces semantic enhancements, ensuring self-containment by resolving dependencies, optimizing algorithms, and transforming trivial snippets into instructive examples, which are absent in Jain et al.’s work (see Section 3.3.2). Although Jain et al. target instruction-tuning data, a smaller and more curated dataset, SwallowCode systematically rewrites large-scale pre-training corpora. This is a more challenging task due to the diversity and volume of code samples. Our pipeline integrates rigorous preprocessing with syntax error and linter-based filtering to ensure high-quality inputs for rewriting.
Taken together, SwallowCode’s transform-and-retain paradigm refines low-quality code rather than discarding it, addressing limitations of existing datasets such as The Stack v1/v2 and Stack-Edu, which rely on filtering. By combining filtering, SGCR, and SCOR, SwallowCode produces a high-quality corpus that significantly improves performance on HumanEval and HumanEval+, advancing the state-of-the-art in open code datasets. Similarly, adhering to the same transform-and-retrain paradigm, SwallowMath rewrites noisy mathematical data, significantly improving performance on GSM8K and MATH.
@@ 2.3 Synthetic Data Generation for Code
Recent approaches, such as Magpie (Xu et al., 2024), leverage high-performance LLMs to generate synthetic instruction-tuning datasets from prompts and desired characteristics without relying on existing data. However, we opted against generating synthetic code data from scratch for SwallowCode due to two key limitations. First, previous work (Chen et al., 2024) demonstrates that low-diversity synthetic data restricts LLM performance. Second, achieving high diversity in synthetic code datasets requires diverse topics and keywords as seeds, as seen in Nemotron-4 340B’s synthetic instruction data (Nvidia et al., 2024). For code, defining such seeds (e.g., varied algorithmic paradigms or problem domains) remains an open challenge, as no established methodology ensures sufficient diversity across programming tasks. Instead, our approach leverages high-quality code from The-Stack-v2, filtered for syntactic and stylistic rigor (Section 3.2), and applies LLM-driven rewriting to enhance quality while preserving the inherent diversity of real-world code. This transform-and-retain strategy maximizes data utility and avoids the risks of synthetic data homogeneity.
@@ 3 Construction of the code corpus
The development of SwallowCode is driven by an empirical and exploratory approach, informed by data ablation experiments evaluating each stage of the data processing pipeline, as illustrated in Figure 2. Specifically, we continually pre-trained the baseline model with the dataset, only differing in the code text subset: before and after applying a specific stage in the pipeline, and decided if we should adopt or discard the stage based on the evaluation results. In this section, we present the experimental results and describe the design choices shaping the pipeline. Please refer to Appendix J.1 for detailed results.
@@ 3.1 Experimental setup
To evaluate the impact of each design choice in our pre-training data processing pipeline, we conduct systematic data ablation studies. Each ablation trains a model that differs only in the target pre-training dataset, holding all other factors constant, including model architecture, parameter count, non-target data, total token budget, and hyperparameters. Specifically, we perform continual pre-training starting from Llama-3.1-8B, using a total of approximately 50 billion tokens. The target dataset is processed with less than one epoch within each ablation. We evaluated model checkpoints approximately every 10 billion tokens using ten downstream benchmarks.
Continual pre-training is conducted using Megatron-LM (Shoeybi et al., 2020). For evaluation, we use evalplus (Liu et al., 2023) and lm-evaluation-harness (Gao et al., 2024) on a suite of benchmarks, with individual tasks detailed in Appendix J. The effectiveness of the code corpus is specifically assessed using HumanEval and HumanEval+. The pre-training data mixture consists of 84% multilingual text and 16% code. Detailed proportions and data source are provided in Appendix A.4.1, with detailed training hyperparameters reported in Appendix A.1. All ablation models, associated checkpoints, and supporting materials are publicly available in our Hugging Face repository  [^1] https://huggingface.co/collections/tokyotech-llm/swallowcode, https://huggingface.co/collections/tokyotech-llm/swallowmath to ensure full reproducibility.
@@ 3.2 Filtering
To construct the SwallowCode corpus, a high-quality Python code dataset, we implement a rigorous filtering pipeline starting with the-stack-v2-train-smol-ids (Lozhkov et al., 2024) as the base dataset. The filtering process is critical to ensure that only syntactically correct and well-structured code is retained, enhancing downstream performance in code generation tasks. We focus exclusively on Python code to maintain consistency across ablation studies and enable fair comparisons with existing public corpora. Our pipeline employs two key filtering techniques, syntax error filtering and linter-based filtering, that significantly improve code quality. We also evaluated LLM-based scoring in ablation experiments, but did not adopt it in the final pipeline due to its limited performance gains relative to computational cost. Figure 3 summarizes the performance of these methods on the HumanEval and HumanEval+ benchmarks, demonstrating the impact of our filtering strategy.
@@ 3.2.1 Syntax error filtering
Despite the heuristic filtering in the BigCode project, the-stack-v2-train-smol-ids includes Python code samples with invalid syntax according to Python 3.10 specifications. To address this, we apply syntax error filtering by compiling each code sample using Python’s built-in compile() function, discarding any samples that fail to compile. This process reduces the dataset from approximately 41 million to 37 million samples, a 9.7% reduction. As shown in Figure 3, removing syntactically invalid samples improves the performance of HumanEval and HumanEval+ [^2] The performance trajectory of The Stack v2 exhibits an initial decline followed by a recovery, consistent with forgetting and subsequent adaptation observed in prior continual pre-training studies (Fujii et al., 2024). This behavior is not particularly notable.. Consequently, we adopt syntax error filtering as a standard step in all subsequent experiments.
@@ 3.2.2 Linter-based filtering
Beyond syntactic correctness, code quality depends on adherence to coding standards. Many samples in the initial dataset exhibit poor structure, generating numerous warnings when analyzed by static analysis tools. We employ pylint, a widely-used Python linter, to enforce a quality threshold, excluding samples with scores below 7.0 on a 0–10 scale based on rule violations. Additionally, we penalize overly verbose comments using a custom heuristic scoring algorithm (detailed in Appendix C). This step reduces the dataset from 36.7 million to 24.1 million samples (34.3% reduction). Figure 3 illustrates the performance gains of more than 1 point in HumanEval and HumanEval+ achieved by linter-based filtering. We first remove syntax errors to avoid wasting computation on files that cannot be linted [^3] While pylint can also surface syntax errors, performing linter-only filtering would be computationally inefficient compared to a staged pipeline of (i) syntax error filtering and (ii) linter-based filtering; the resulting retained set is effectively the same up to linter version/config nuances. . Consequently, we adopt linter-based filtering, with a threshold of 7.0, as a standard step in all subsequent experiments.
3.2.3 LLM-based score filtering
Recent approaches leverage LLM to generate synthetic annotations for training quality classifiers, which are then used to filter web-scale corpora by retaining high-quality samples (e.g., FineWeb (Penedo et al., 2024), Stack-Edu). Instead of training a separate classifier, we directly prompt Llama-3.3-70B-Instruct to evaluate each Python code snippet on a scale of 0–10, based on ten criteria, including code readability, modularity, and adherence to naming conventions, derived from the Google Python Style Guide. The detailed scoring prompt and the distribution of the quality scores are provided in Appendix D.